# A Minimal Quantitative Model of Perceptual Suppression and Breakthrough in Visual Rivalry

### v1

@@ 5 Confirmation of novel prediction
Crucially, the symmetry of dominance and suppression durations is a prediction that we can test in the empirical data of Alais and colleagues [47]. If we assume that the empirical data are generated by a noisy version of an equivalent dynamical system, the distribution of dominance durations should be approximately equal to the distribution of suppression durations. To test this prediction, we took a two-pronged approach. First, we aggregated the suppression and dominance duration data across all three experiments from Alais and colleagues [47]. We excluded trials with dominance or suppression durations longer than the time it takes for the stimulus to move from the minimum to the maximum allowed contrast value, as well as individual periods of dominance or suppression shorter than 500ms, as these instances likely reflect either a noise-driven perceptual switch or a failure to follow task instructions. In addition, as the prediction applies to systems close to their equilibrium value, we excluded the first period of dominance and suppression after trial onset. We then used a two-sample Kolmogorov-Smirnov test against the null hypothesis that the data come from the same underlying distribution. The test failed to reject the null hypothesis at the 5% significance level ($p=0.093$).
Second, having failed to reject the null hypothesis that the data come from the same underlying distribution, we next sought to find confirmatory evidence for the equivalence of the dominance and suppression distributions. We therefore fit gamma and lognormal distributions, which have both previously been shown to provide a good fit to bistable perception data [5, 52, 6, 53, 27, 42], to the dominance and suppression duration data via maximum likelihood estimation. Both distributions provided an excellent fit to the data (Figure 3D-F) with the gamma distribution performing marginally better than the lognormal distribution in terms of the negative log likelihoods (dominant $l_{\text{gamma}}/l_{\text{lognormal}}=0.9933$, suppressed $l_{\text{gamma}}/l_{\text{lognormal}}=0.9977$)[[, but, also see,]]Brascamp2005. In line with the prediction from the model, parameter estimates for dominance and suppression distributions were near identical for both the gamma ($\alpha_{\text{Dom}}=4.54,\theta_{\text{Dom}}=0.79$; $\alpha_{\text{Sup}}=4.50,\theta_{\text{Sup}}=0.81$) and lognormal ($\mu_{\text{Dom}}=1.18,\sigma_{\text{Dom}}=0.49$; $\mu_{\text{Sup}}=1.16,\sigma_{\text{Sup}}=0.48$) distributions with almost entirely overlapping 95% confidence intervals (Figure 3F).
@@ 7 Discussion
$\tau_{E}$ Neuronal population time constant (ms) 15 ms
$\tau_{H}$ Hyperpolarizing adaptation current time constant (ms) 1000 ms
$M$ Constant drive of mask (a.u.) 0.8 (a.u.)
$S$ Initial condition for stimulus drive $S_{t_{0}}=1.2$ (a.u.)
$\varepsilon$ Strength of excitatory recurrent projections 0.05 (a.u.)
$a$ Strength of competitive inhibition 3.3-3.5 (base = 3.4) (a.u.)
$g_{M}$ Strength of mask population adaptation current 1.7 (a.u.)
$g_{S}$ Strength of stimulus population adaptation current 3 (a.u.)
$\gamma$ Rate of contrast change ($0.0021-0.0063)\times 10^{-2}$ (a.u.)
@@
Starting with line one of equation 9, where for ease of notation we let $E_{M}(\infty)=M/(1+g_{M}-\varepsilon)$ and $E_{S}(\infty)=S_{S}/(1+g_{S}-\varepsilon)$, we have:
$S_{S}+\gamma t-aE_{M}(\infty)-g_{S}E_{S}(\infty)e^{-t/\tau_{H}}$ $=0$
$aE_{M}(\infty)-S_{S}-\gamma t$ $=-g_{S}E_{S}(\infty)e^{-t/\tau_{H}}$
$(aE_{M}(\infty)-S_{S}-\gamma t)e^{t/\tau_{H}}$ $=-g_{S}E_{S}(\infty)$
$\left(t+\frac{S_{S}-aE_{M}(\infty)}{\gamma}\right)e^{t/\tau_{H}}$ $=\frac{g_{S}E_{S}(\infty)}{\gamma}$
$\left(\frac{t}{\tau_{H}}+\frac{S_{S}-aE_{M}(\infty)}{\gamma\tau_{H}}\right)e^{\frac{t}{\tau_{H}}+\frac{S_{S}-aE_{M}(\infty)}{\gamma\tau_{H}}}$ $=\frac{g_{S}E_{S}(\infty)}{\gamma\tau_{H}}e^{\frac{S_{S}-aE_{M}(\infty)}{\gamma\tau_{H}}}$
$\frac{t}{\tau_{H}}+\frac{S_{S}-aE_{M}(\infty)}{\gamma\tau_{H}}$ $=W_{0}\left(\frac{g_{S}E_{S}(\infty)}{\gamma\tau_{H}}e^{\frac{S_{S}-aE_{M}(\infty)}{\gamma\tau_{H}}}\right)\,.$
After minor rearrangement and expanding both $E_{M}(\infty)$ and $E_{S}(\infty)$ we obtain,
@@
Correction in hand, we let $E_{S}(S_{B})=\tilde{S}_{B}/(1+g_{S}-\varepsilon)$, to arrive at,
$M-\frac{a(\tilde{S}_{B}-\gamma t)}{1+g_{S}-\varepsilon}-g_{M}E_{M}(\infty)e^{-t/\tau_{H}}$ $=0$
$M-\frac{a\tilde{S}_{B}}{1+g_{S}-\varepsilon}+\frac{a\gamma t}{1+g_{S}-\varepsilon}$ $=g_{M}E_{M}(\infty)e^{-t/\tau_{H}}$
$\left(M-\frac{a\tilde{S}_{B}}{1+g_{S}-\varepsilon}+\frac{a\gamma t}{1+g_{S}-\varepsilon}\right)e^{t/\tau_{H}}$ $=g_{M}E_{M}(\infty)$
$\left(\frac{t}{\tau_{H}}-\frac{a\tilde{S}_{B}-M(1+g_{S}-\varepsilon)}{a\gamma\tau_{H}}\right)e^{\frac{t}{\tau_{H}}-\frac{a\tilde{S}_{B}-M(1+g_{S}-\varepsilon)}{a\gamma\tau_{H}}}$ $=\frac{g_{M}E_{M}(\infty)(1+g_{S}-\varepsilon)}{a\gamma\tau_{H}}e^{-\frac{a\tilde{S}_{B}-M(1+g_{S}-\varepsilon)}{a\gamma\tau_{H}}}$
$\frac{t}{\tau_{H}}-\frac{a\tilde{S}_{B}-M(1+g_{S}-\varepsilon)}{a\gamma\tau_{H}}$ $=W_{0}\left(\frac{g_{M}E_{M}(\infty)(1+g_{S}-\varepsilon)}{a\gamma\tau_{H}}e^{-\frac{a\tilde{S}_{B}-M(1+g_{S}-\varepsilon)}{a\gamma\tau_{H}}}\right)\,$,
expanding $E_{S}(S_{B})$ and solving for $t$, we obtain,

### v2

@@ 5 Confirmation of novel prediction
Crucially, the symmetry of dominance and suppression durations is a prediction that we can test in the empirical data of Alais and colleagues [47]. If we assume that the empirical data are generated by a noisy version of an equivalent dynamical system, the distribution of dominance durations should be approximately equal to the distribution of suppression durations. To test this prediction, we took a two-pronged approach. First, we aggregated the suppression and dominance duration data across all three experiments from Alais and colleagues [47]. We excluded trials with dominance or suppression durations longer than the time it takes for the stimulus to move from the minimum to the maximum allowed contrast value, as well as individual periods of dominance or suppression shorter than 500ms, as these instances likely reflect either a noise-driven perceptual switch or a failure to follow task instructions. In addition, as the prediction applies to systems close to their equilibrium value, we excluded the first two periods of dominance and suppression after trial onset. We then used a two-sample Kolmogorov-Smirnov test against the null hypothesis that the data come from the same underlying distribution. The test failed to reject the null hypothesis at the 5% significance level ($p=0.5555$).
Second, having failed to reject the null hypothesis that the data come from the same underlying distribution, we next sought to find confirmatory evidence for the equivalence of the dominance and suppression distributions. We therefore fit gamma and lognormal distributions, which have both previously been shown to provide a good fit to bistable perception data [5, 52, 6, 53, 27, 42], to the dominance and suppression duration data via maximum likelihood estimation. Both distributions provided an excellent fit to the data (Figure 3D-F) with the gamma distribution performing marginally better than the lognormal distribution in terms of the negative log likelihoods (dominant $l_{\text{gamma}}/l_{\text{lognormal}}=0.9945$, suppressed $l_{\text{gamma}}/l_{\text{lognormal}}=0.9971$)[[, but, also see,]]Brascamp2005. In line with the prediction from the model, parameter estimates for dominance and suppression distributions were near identical for both the gamma ($\alpha_{\text{Dom}}=4.64,\theta_{\text{Dom}}=0.784$; $\alpha_{\text{Sup}}=4.63,\theta_{\text{Sup}}=0.78$) and lognormal ($\mu_{\text{Dom}}=1.18,\sigma_{\text{Dom}}=0.49$; $\mu_{\text{Sup}}=1.17,\sigma_{\text{Sup}}=0.48$) distributions with almost entirely overlapping 95% confidence intervals (Figure 3F).
@@ 7 Discussion
$\tau_{E}$ Neuronal population time constant 15 ms
$\tau_{H}$ Hyperpolarizing adaptation current time constant 1000 ms
$M$ Constant drive of mask 0.8 a.u.
$S$ Initial condition for stimulus drive $S_{t_{0}}=1.2$ a.u.
$\varepsilon$ Strength of excitatory recurrent projections 0.05 a.u.
$a$ Strength of competitive inhibition 3.3-3.5 (base = 3.4) a.u.
$g_{M}$ Strength of mask population adaptation current 1.7 a.u.
$g_{S}$ Strength of stimulus population adaptation current 3 a.u.
$\gamma$ Rate of contrast change ($0.0021-0.0063)\times 10^{-2}$ a.u.
@@
Starting with line one of equation 9, where for compactness of notation we let $E_{M}=E_{M}(\infty)=M/(1+g_{M}-\varepsilon)$ and $E_{S}=E_{S}(\infty)=S_{S}/(1+g_{S}-\varepsilon)$, we have:
$S_{S}+\gamma t-aE_{M}-g_{S}E_{S}e^{-t/\tau_{H}}$ $=0$
$aE_{M}-S_{S}-\gamma t$ $=-g_{S}E_{S}e^{-t/\tau_{H}}$
$(aE_{M}-S_{S}-\gamma t)e^{t/\tau_{H}}$ $=-g_{S}E_{S}$
$\left(t+\frac{S_{S}-aE_{M}}{\gamma}\right)e^{t/\tau_{H}}$ $=\frac{g_{S}E_{S}}{\gamma}$
$\left(\frac{t}{\tau_{H}}+\frac{S_{S}-aE_{M}}{\gamma\tau_{H}}\right)e^{\frac{t}{\tau_{H}}+\frac{S_{S}-aE_{M}}{\gamma\tau_{H}}}$ $=\frac{g_{S}E_{S}}{\gamma\tau_{H}}e^{\frac{S_{S}-aE_{M}}{\gamma\tau_{H}}}$
$\frac{t}{\tau_{H}}+\frac{S_{S}-aE_{M}}{\gamma\tau_{H}}$ $=W_{0}\left(\frac{g_{S}E_{S}}{\gamma\tau_{H}}e^{\frac{S_{S}-aE_{M}}{\gamma\tau_{H}}}\right)\,.$
After minor rearrangement and expanding both $E_{M}$ and $E_{S}$ we obtain,
@@
Correction in hand we arrive at,
$M-\frac{a(\tilde{S}_{B}-\gamma t)}{1+g_{S}-\varepsilon}-g_{M}E_{M}e^{-t/\tau_{H}}$ $=0$
$M-\frac{a\tilde{S}_{B}}{1+g_{S}-\varepsilon}+\frac{a\gamma t}{1+g_{S}-\varepsilon}$ $=g_{M}E_{M}e^{-t/\tau_{H}}$
$\left(M-\frac{a\tilde{S}_{B}}{1+g_{S}-\varepsilon}+\frac{a\gamma t}{1+g_{S}-\varepsilon}\right)e^{t/\tau_{H}}$ $=g_{M}E_{M}$
$\left(\frac{t}{\tau_{H}}-\frac{a\tilde{S}_{B}-M(1+g_{S}-\varepsilon)}{a\gamma\tau_{H}}\right)e^{\frac{t}{\tau_{H}}-\frac{a\tilde{S}_{B}-M(1+g_{S}-\varepsilon)}{a\gamma\tau_{H}}}$ $=\frac{g_{M}E_{M}(1+g_{S}-\varepsilon)}{a\gamma\tau_{H}}e^{-\frac{a\tilde{S}_{B}-M(1+g_{S}-\varepsilon)}{a\gamma\tau_{H}}}$
$\frac{t}{\tau_{H}}-\frac{a\tilde{S}_{B}-M(1+g_{S}-\varepsilon)}{a\gamma\tau_{H}}$ $=W_{0}\left(\frac{g_{M}E_{M}(1+g_{S}-\varepsilon)}{a\gamma\tau_{H}}e^{-\frac{a\tilde{S}_{B}-M(1+g_{S}-\varepsilon)}{a\gamma\tau_{H}}}\right)\,$,
expanding $E_{M}$ and solving for $t$, we obtain,